OpenStudy (hi54321):

first to give me answer gets faned and gets a medal What is the slope-intercept form of the function that contains the points (6, 2) and (4, 8)? y = x + @anyone123

OpenStudy (hi54321):

@amistre64 @Nnesha @whydoihavetosignup1 @rational @sleepyhead314 @hhelpplzzzz @andrewhaze @Mesopretty @Rizags

OpenStudy (sleepyhead314):

first find the slope \[slope = \frac{ rise }{ run } = \frac{ y _{2}-y _{1} }{ x _{2}-x _{1} } \]when given points (x1, y1) and (x2, y2)

OpenStudy (sleepyhead314):

so you will get 8 - 2 slope = ------ = ? 4 - 6

OpenStudy (hi54321):

so is that the answer @sleepyhead314 because I really suck at this

OpenStudy (sleepyhead314):

not quite :) I'm trying to help you step by step 8 - 2 slope = ------ = ? 4 - 6 try to simplify that :)

OpenStudy (hi54321):

so 6 and -2

OpenStudy (jdoe0001):

why not just solve it first? that way you'd get the medal

OpenStudy (sleepyhead314):

right then 6 divided by (-2) = ?

OpenStudy (hi54321):

-3

OpenStudy (sleepyhead314):

good so now you have your slope ^_^ slope = -3 then we have to find our y-intercept

OpenStudy (sleepyhead314):

y = -3x + b we can temporarily plug in the point (6, 2) for (x, y) into the equation to solve for b 2 = (-3)*(6) + b can you solve for b from there? :)

OpenStudy (hi54321):

I don't know? are we almost to the answer though @sleepyhead314

OpenStudy (sleepyhead314):

yes, all you have to do is solve for b from the equation 2 = (-3)*(6) + b I'll do a few steps for you 2 = -18 + b then add 18 to both sides

OpenStudy (hi54321):

so 20 = -18 + 18b?

OpenStudy (sleepyhead314):

whoops you're thinking too hard ^_^ b = 2 + 18 = ?

OpenStudy (hi54321):

oh so just 20?

OpenStudy (sleepyhead314):

yes :) then we plug in the numbers we got back into the slope intercept form! y = m x + b m = -3 b = 20 just plug in those two numbers!

OpenStudy (hi54321):

17?

OpenStudy (sleepyhead314):

you don't need to add or subtract anything :) just PLUG the two numbers for the Letters y = (m) x + (b) m = -3 b = 20

OpenStudy (hi54321):

oh ok thanks :)